# DeRIS: Decoupling Perception and Cognition for Enhanced Referring Image Segmentation through Loopback Synergy

### v1

@@ Abstract
Referring Image Segmentation (RIS) is a challenging task that aims to segment objects in an image based on natural language expressions. While prior studies have predominantly concentrated on improving vision-language interactions and achieving fine-grained localization, a systematic analysis of the fundamental bottlenecks in existing RIS frameworks remains underexplored. To bridge this gap, we propose DeRIS, a novel framework that decomposes RIS into two key components: perception and cognition. This modular decomposition facilitates a systematic analysis of the primary bottlenecks impeding RIS performance. Our findings reveal that the predominant limitation lies not in perceptual deficiencies, but in the insufficient multi-modal cognitive capacity of current models. To mitigate this, we propose a Loopback Synergy mechanism, which enhances the synergy between the perception and cognition modules, thereby enabling precise segmentation while simultaneously improving robust image-text comprehension. Additionally, we analyze and introduce a simple non-referent sample conversion data augmentation to address the long-tail distribution issue related to target existence judgement in general scenarios. Notably, DeRIS demonstrates inherent adaptability to both non- and multi-referents scenarios without requiring specialized architectural modifications, enhancing its general applicability. The codes and models are available at https://github.com/Dmmm1997/DeRIS
@@ 2.1 Referring Image Segmentation (RIS)
Perception-centric RIS methods are typically categorized into two structural paradigms: post-fusion and early-fusion. Post-fusion methods [53, 39, 31, 50, 44, 82, 48, 27] primarily focus on the fusion of visual and linguistic features through specially designed interaction modules that are applied after the visual backbone. In contrast, early-fusion methods [15, 19, 73, 46, 74, 80, 10, 63, 5] emphasize integrating visual and linguistic features during the feature extraction phase via modularized architectural designs, facilitating deeper cross-modal understanding. Cognition-centric RIS methods [11, 77, 71, 12] leverage powerful pre-trained multimodal models [57, 1, 32, 66] to generate aligned or fused multimodal representations, thereby enhancing the model’s cognitive capacity for interpreting complex image-text relationships. In this work, we introduce a novel architectural framework that decouples the perception and cognition components, effectively harnessing the strengths of both paradigms to advance RIS performance.
@@ 3.2.2 Perception Layer
where Conv denotes a 1$\times$1 convolution operation. This design primarily aims to bolster the model’s capability to generate text-informed proposals by incorporating contextual semantic information from the cognition branch, while simultaneously enhancing its perceptual accuracy.
@@ 3.4 Training Objectives
$\mathcal{L}^{i}=\lambda_{m}\mathcal{L}_{mask}^{i}+\lambda_{r}\mathcal{L}_{r}^{i}+\lambda_{nt}\mathcal{L}_{nt}^{i},\vspace{-5pt}$ (7)
$\mathcal{L}=\lambda_{aux}\sum_{i=1}^{N_{r}-1}{\mathcal{L}^{i}}+\mathcal{L}^{N_{r}},\vspace{-5pt}$ (8)
@@
ReLA [CVPR’23] [45] Swin-B BERT-B 73.82 76.48 70.18 66.04 71.02 57.65 65.00 65.97
LAVT [CVPR’22] [73] Swin-B BERT-B 74.46 76.89 70.94 65.81 70.97 59.23 63.34 63.62
@@ 4.1 Settings
We conduct extensive experiments on four publicly available datasets: RefCOCO/+/g [75, 55], gRefCOCO [45]. For the perception branch, we initialize the model using pretrained weights from Mask2Former [9] and use an input resolution of 384$\times$384. For the cognition branch, we initialize the model using pretrained weights from BEiT3 [66] and use an input resolution of 224$\times$224. Additional implementation details are provided in the Appendix C.
@@ 4.4.2 Impact of Conversion Rate of Non-referent
The proposed NSC data augmentation strategy transforms image-sentence pairs from referent to non-referent cases. The conversion probability $R_{c}$ serves as a critical hyperparameter in this process. As shown in Table 8, we observe that increasing $R_{c}$ leads to a steady improvement in non-referent accuracy (N-acc). However, when $R_{c}$ exceeds a certain threshold, performance on other metrics, such as cIoU, begins to decline. This occurs primarily because enhancing the diversity of non-referent samples prompts the model to assign higher confidence to predicting the absence of a target, effectively mitigating the low-confidence issue in tail categories stemming from the long-tail distribution of the dataset. In this paper, we set $R_{c}$=15to address the limited diversity and scarcity of non-referent samples.
@@ Acknowledgements
This work is supported by the National Natural Science Foundation of China under Nos. 62276061 and 62436002. This work is also supported by Research Fund for Advanced Ocean Institute of Southeast University (Major Program MP202404). This work is also supported by the SEU Innovation Capability Enhancement Plan for Doctoral Students.
Supplementary Material
Appendix A Datasets
A.1 RefCOCO Series
The RefCOCO series datasets [75, 55] constitute the primary benchmarks for Referring Image Segmentation (RIS):
  • RefCOCO contains 142,209 referring expressions for 50,000 objects across 19,994 images. It features two test splits: testA (focus on images with multiple people) and testB (focus on images with multiple object instances).
  • RefCOCO+ comprises 141,564 expressions with stricter annotation protocols that prohibit explicit spatial relationships, making it more challenging than RefCOCO.
  • RefCOCOg contains 85,474 linguistically complex expressions (average length 8.4 words) describing 54,822 objects in 26,711 images.
The results presented in this paper are based on the Mixed dataset, which combines the RefCOCO [75], RefCOCO+ [75], and RefCOCOg-umd [55] datasets.
A.2 gRefCOCO
The gRefCOCO dataset [23, 45] extends traditional RIS with two critical enhancements:
  • Multi-target references: 80,022 expressions describing multiple objects
  • Non-referent expressions: 32,202 descriptions without corresponding targets
This comprehensive benchmark contains 278,232 expressions across 60,287 object instances in 19,994 images, following the standard RefCOCO [75] UNC partition scheme (train/val/testA/testB).
Appendix B Evaluation Metrics
For RIS, we use mean Intersection over Union (mIoU) as the evaluation metric. For GRES, we evaluate our model using three metrics: gIoU, cIoU, and N-acc on gRefCOCO [45]. The gIoU calculates the average IoU across all instances in an image, treating empty targets as true positives with an IoU of 1. The cIoU evaluates the ratio of intersection to union pixels. N-acc in gRefCOCO represents the proportion of correctly classified non-referent expressions.
Appendix C Additional Implementation Details
The model is trained using the Adam [33] optimizer with a learning rate of $1\times 10^{-4}$ and a batch size of 32. For the GRES task, the maximum sentence length is limited to 50 words, and the number of object queries is set to 20. The model is trained for 6 epochs, with the learning rate reduced by a factor of 0.1 at the 5th epoch. For the RIS task, the maximum sentence length is limited to 20 words, and the number of object queries is set to 10. The model is trained for 20 epochs, with the learning rate reduced by a factor of 0.1 at the 15th and 19th epoch. The experimental results are conducted in 8 NVIDIA A100(40G) GPUs. Ablation studies are conducted on the val set of gRefCOCO [45] without additional explanations.
For the large language model version, we employ LoRA [24] for efficient training, setting the rank to 8 and the alpha to 16. We train the model for 10 epochs on a mixed dataset comprising the RefCOCO series. The training batch size per GPU is set to 4, and we use a learning rate of 4e-5, with AdamW [52] as the optimizer.
Appendix D Additional Methodology
D.1 Text Similarity Metrics
Jaccard similarity is a set-based similarity measure that focuses on the presence of key terms, making it well-suited for comparing short texts. Given two text segments represented as word sets $A$ and $B$, the Jaccard similarity is defined as:
$J(A,B)=\frac{|A\cap B|}{|A\cup B|}$ (9)
where $A\cap B$ denotes the intersection of the word sets, and $A\cup B$ represents their union.
Cosine similarity, on the other hand, is a vector-based measure that incorporates word frequency weights, making it more effective for longer texts. For vectorized text representations $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$, the cosine similarity is computed as:
$\cos(\theta)=\frac{\mathbf{v}_{1}\cdot\mathbf{v}_{2}}{\|\mathbf{v}_{1}\|\|\mathbf{v}_{2}\|}$ (10)
where $\cdot$ denotes the dot product, and $\|\mathbf{v}\|$ represents the Euclidean norm of the vector. To compute text similarity, we evaluate both set-based and vector-based similarities. The final similarity score is obtained by linearly combining the two measures:
$S_{\text{final}}=\frac{1}{2}\left(J(A,B)+\cos(\theta)\right)$ (11)
Through experiments, we determine an optimal threshold $T_{s}$, above which two text segments are considered similar.
Appendix E Additional Experimental Results
E.1 Impact of the Number of Loopback Synergy Rounds
Figure 7: Impact of the number of loopback synergy rounds $N_{r}$ on both training efficiency and performance. We ultimately select $N_{r}=3$ to balance efficiency and effectiveness.
The multi-round interaction mechanism of loopback synergy, along with the auxiliary loss computation, introduces additional computational complexity during both forward and backward propagation. To evaluate the impact of different interaction rounds $N_{r}$ on model performance and training efficiency, we conduct experiments. As shown in Fig. 7, we observe that increasing $N_{r}$ leads to performance improvements, but at the cost of reduced training efficiency. Based on this trade-off, we select $N_{r}=3$ as a compromise between performance and efficiency.
E.2 Impact of the Number of Object Queries
Figure 8: Impact of the number of object queries $N_{q}$ on model performance.
The effect of the number of object queries $N_{q}$ on model performance is shown in Fig. 8. Given that the number of referents in the gRefCOCO [45] dataset typically does not exceed 10, we observe that increasing $N_{q}$ leads to a decline in model performance. This performance degradation primarily arises because a larger number of object queries increases the probability of similar masks being assigned to each query, which in turn amplifies the ambiguity in the one-to-one matching process and makes the referent classification task more challenging. This additional complexity indirectly slows down convergence. Consequently, we select $N_{q}=20$ as a balanced choice to optimize both performance and efficiency.
E.3 Impact of the Source of Non-referent Judgement
Table 9: The impact of source on non-referent judgement.
Source N-acc. gIoU cIoU MaskAcc@-0.9
query feat 70.54 70.24 66.30 21.03
image feat 69.65 69.45 65.89 20.13
cls feat 69.78 69.54 65.75 20.10
The non-referent is designed to determine whether the target referred to by the expression is present in the image, which directly impacts the model’s performance. we conduct experiments to investigate the influence of different sources of non-referent on performance. As shown in Table 9, the sources for non-referent judgment, denoted as query, image, and cls, correspond to the query from loopback synergy, the image features and the global features from the image-text pre-trained encoder, respectively. The results indicate that query significantly outperforms the other two alternatives in terms of non-referent discrimination performance. We attribute this superior performance to the fact that the cognition branch alone lacks the fine-grained spatial perception capabilities. In contrast, the query features, through multiple rounds of loopback synergy, progressively enhance the discrimination of non-referent.
Table 10: The impact of weight-sharing in cognition decoder.
Strategy N-acc. gIoU cIoU MaskAcc@-0.9
Share 67.83 68.65 65.50 19.58
Independent 70.54 70.24 66.30 21.03
Figure 9: Visualization results of CRIS, LAVT, and DeRIS on the RefCOCOg test set.
E.4 Impact of the Weight Sharing in Loopback Synergy
The weights in the cognition decoder can either be shared or independent. we conducted experiments to compare the impact of weight-sharing and weight-independent of the cognition decoder on the performance. As shown in Table 10, the results demonstrate that the weight-independent strategy significantly outperforms the weight-sharing strategy in terms of performance. We attribute this to the fact that independent weights allow for a more precise representation of features at each round, mitigating the representation coupling issue caused by weight-sharing. This, in turn, facilitates more effective model training and optimization.
Figure 10: Visualization of predicted masks and instance masks with confidence $>$ 0.5 on the gRefCOCO dataset.
Figure 11: Visualization comparison of the NSC strategy in non-referent target scenarios.
Figure 12: Visualization comparison of the NSC strategy in referent target scenarios.
Table 11: Loopback Rounds Efficiency Analysis. Train time is reported on GRefCOCO.
cIoU MACs Param(M) Infer(ms) Train(min)
1 64.9 37.0 230.9 22.5 146.8
3 66.4 37.1 235.9 23.4 160.6
5 66.7 37.2 240.9 23.9 182.4
E.5 Loopback Efficiency Analysis
As shown in the Table 11, increasing the recurrence count from 1 to 3 leads to a 1.5% improvement in the cIoU metric, demonstrating the necessity of deeper interaction between the perception and understanding branches. However, further increasing the count to 5 yields only a marginal gain of 0.3% in cIoU, while significantly increasing both model parameters and training time. To balance performance and efficiency, we set the recurrence count to 3.
Table 12: Loopback Rounds Efficiency Analysis. Train time is reported on GRefCOCO.
Method Ref-YTB MeVIS(valid-u)
J F J&F J F J&F
DeRIS-B 58.4 59.6 59.0 44.7 51.8 48.2
DeRIS-L 62.7 67.4 65.1 48.8 55.9 52.4
E.6 Performance on Video-Specific Datasets
As shown in Table 12, we report zero-shot performance on the Ref-YouTube-VOS [60] and MeVIS [16] datasets. Despite not utilizing temporal information, our DeRIS framework still achieves competitive results, demonstrating its strong potential for extension to video grounding scenarios.
Appendix F Qualitative Analyses
F.1 Comparison with Different Centric Architecture Models
Fig. 9 presents the visualization results of CRIS, LAVT, and DeRIS on the RefCOCOg test set. The results demonstrate that the DeRIS model effectively segments the referent targets in complex scenes, producing high-quality instance masks with clear boundaries and accurate segmentation. Moreover, the model successfully captures the fine-grained details of the referent targets. LAVT adopts a perception-centric architecture, exhibiting strong fine-grained perception capabilities. However, it struggles with comprehensive vision-language understanding. In contrast, CRIS (cognition-centric) excels in precise vision-language comprehension but lacks fine-grained perception, often resulting in coarse and imprecise contours. DeRIS integrates the strengths of both architectures, achieving a robust balance between fine-grained perception and semantic understanding.
F.2 Visualization of Instance Masks on gRefCOCO
Fig. 10 presents the visualization of instance masks with confidence scores greater than 0.5 on the gRefCOCO dataset. It can be observed that even in complex multi-object scenes, DeRIS accurately segments the target instances while ensuring strong textual relevance. This is attributed to its robust perceptual capabilities as a prerequisite, coupled with a powerful cognitive model for context understanding. Furthermore, even in challenging scenarios involving extremely small objects, such as the fourth sample, DeRIS remains capable of precisely capturing the target and assigning a high confidence score.
Figure 13: Visualization of segmentation results under different textual conditions on the gRefCOCO dataset.
F.3 Visualization Analysis of the NSC Strategy
F.3.1 Non-Referent Target Scenarios
Fig. 11 illustrates the effectiveness of the NSC strategy in handling non-referent samples. Taking (a) as an example, the results demonstrate that without the NSC strategy, the model erroneously identifies the woman on the right, wearing a mixed-color dress, as the referent target with a confidence score of 1.00, leading to a clear misclassification. However, after applying the NSC strategy, the model correctly classifies this instance as a non-referent target, with its confidence score significantly reduced to 0.04, indicating a substantial improvement. This comparative experiment validates that the NSC strategy enhances the model’s ability to distinguish between visually similar distracting objects, thereby improving its robustness in recognizing non-referent instances within complex scenes.
Figure 14: Visualization of RefCOCO/+/g dataset.
F.3.2 Referent Existing Target Scenarios
Fig. 12 presents a performance comparison of the NSC strategy in scenarios involving referent targets. Taking (a) as an example, the results show that the ”velvet devil” pattern on the label of the second bottle and the floral pattern on the label of the third bottle are valid referent targets, while the other two bottles act as distracting instances. In the baseline model, without the NSC strategy, two common errors occur: false positives for the non-referent bottles (with confidence scores of 0.96 and 0.93), and a missed detection of the floral-patterned bottle, which has a confidence score of 0.64—below the detection threshold of 0.7.
After incorporating the NSC strategy during training, the model exhibits significant improvements: the confidence scores of the distracting samples decrease to 0.63 and 0.02, while the confidence score of the floral-patterned bottle rises to 0.96, successfully ensuring accurate detection. These results indicate that the NSC strategy enhances the model’s ability to suppress false positives in complex scenes by establishing more differentiated representations between targets while significantly improving its detection capability for weakly salient referent targets.
F.4 Visualization of Different Textual Conditions on gRefCOCO
Fig. 13 presents the segmentation results under different textual conditions on the gRefCOCO dataset. Through visualizing the model’s response to various textual expressions for the same image, we observe that DeRIS demonstrates strong capabilities in comprehending both text and images, as well as effectively identifying cases where the referent object is absent. Even in scenarios involving long and complex textual descriptions, visually intricate targets, and challenging cases with no referent object, the model exhibits remarkable robustness and discrimination ability.
F.5 Visualization of RefCOCO/+/g Dataset
Fig. 14 presents the visualization results on the RefCOCO/+/g dataset. In one-to-one correspondence scenarios, DeRIS also demonstrates strong performance without requiring additional architectural modifications. The model seamlessly adapts to single-object, multi-object, and absent-object scenarios, maintaining robust and consistent performance across diverse cases.

### v2

# DeRIS: Decoupling Perception and Cognition for Enhanced Enhanced Referring Image Segmentation through Loopback Synergy

@@ Abstract
Referring Image Segmentation (RIS) is a challenging task that aims to segment objects in an image based on natural language expressions. While prior studies have predominantly concentrated on improving vision-language interactions and achieving fine-grained localization, a systematic analysis of the fundamental bottlenecks in existing RIS frameworks remains underexplored. To bridge this gap, we propose DeRIS, a novel framework that decomposes RIS into two key components: perception and cognition. This modular decomposition facilitates a systematic analysis of the primary bottlenecks impeding RIS performance. Our findings reveal that the predominant limitation lies not in perceptual deficiencies, but in the insufficient multi-modal cognitive capacity of current models. To mitigate this, we propose a Loopback Synergy mechanism, which enhances the synergy between the perception and cognition modules, thereby enabling precise segmentation while simultaneously improving robust image-text comprehension. Additionally, we analyze and introduce a simple non-referent sample conversion data augmentation to address the long-tail distribution issue related to target existence judgement in general scenarios. Notably, DeRIS demonstrates inherent adaptability to both non- and multi-referents scenarios without requiring specialized architectural modifications, enhancing its general applicability. The codes and models are available at link
@@ 2.1 Referring Image Segmentation (RIS)
Perception-centric RIS methods are typically categorized into two structural paradigms: post-fusion and early-fusion. Post-fusion methods [51, 38, 31, 49, 43, 79, 47, 27] primarily focus on the fusion of visual and linguistic features through specially designed interaction modules that are applied after the visual backbone. In contrast, early-fusion methods [17, 20, 70, 45, 71, 77, 10, 60, 5] emphasize integrating visual and linguistic features during the feature extraction phase via modularized architectural designs, facilitating deeper cross-modal understanding. Cognition-centric RIS methods [11, 74, 68, 14, 12, 13] leverage powerful pre-trained multimodal models [55, 1, 32, 63] to generate aligned or fused multimodal representations, thereby enhancing the model’s cognitive capacity for interpreting complex image-text relationships. In this work, we introduce a novel architectural framework that decouples the perception and cognition components, effectively harnessing the strengths of both paradigms to advance RIS performance.
@@ 3.2.2 Perception Layer
where Conv denotes a $\times1$ convolution operation. This design primarily aims to bolster the model’s capability to generate text-informed proposals by incorporating contextual semantic information from the cognition branch, while simultaneously enhancing its perceptual accuracy.
@@ 3.4 Training Objectives
$\mathcal{L}^{i}=\lambda_{m}\mathcal{L}_{mask}^{i}+\lambda_{r}\mathcal{L}_{r}^{i}+\lambda_{nt}\mathcal{L}_{nt}^{i},\vskip-5.0pt$ (7)
$\mathcal{L}=\lambda_{aux}\sum_{i=1}^{N_{r}-1}{\mathcal{L}^{i}}+\mathcal{L}^{N_{r}},\vskip-5.0pt$ (8)
@@ 4.1 Settings
We conduct extensive experiments on four publicly available datasets: RefCOCO/+/g [72, 53], gRefCOCO [44]. For the perception branch, we initialize the model using pretrained weights from Mask2Former [9] and use an input resolution of $\times384$. For the cognition branch, we initialize the model using pretrained weights from BEiT3 [63] and use an input resolution of $\times224$. Additional implementation details are provided in the Appendix C.
@@ 4.4.2 Impact of Conversion Rate of Non-referent
The proposed NSC data augmentation strategy transforms image-sentence pairs from referent to non-referent cases. The conversion probability $R_{c}$ serves as a critical hyperparameter in this process. As shown in Table 8, we observe that increasing $R_{c}$ leads to a steady improvement in non-referent accuracy (N-acc). However, when $R_{c}$ exceeds a certain threshold, performance on other metrics, such as cIoU, begins to decline. This occurs primarily because enhancing the diversity of non-referent samples prompts the model to assign higher confidence to predicting the absence of a target, effectively mitigating the low-confidence issue in tail categories stemming from the long-tail distribution of the dataset. In this paper, we set $R_{c}$=15% to address the limited diversity and scarcity of non-referent samples.
@@ Acknowledgements
This work is supported by the National Natural Science Foundation of China under Nos. 62276061 and 62436002. This work is also supported by Research Fund for Advanced Ocean Institute of Southeast University (Major Program MP202404). This work is also supported by the SEU Innovation Capability Enhancement Plan for Doctoral Students (CXJH_SEU25125).